Please explain to me the proper way of solving this problem.
Could you please include the answer also?
1) Christina has $22.00 more than 3 times as much money as Melissa has. Together they have $122.00. How much money does each girl have?
I need to write the equation :
Solve for "n" :
Answer:
Thanks..

Answer by stanbon(75887)

1) Christina has $22.00 more than 3 times as much money as Melissa has. Together they have $122.00. How much money does each girl have?
Let amount Melissa has be "n".
Then amount Christina has is "3n+22"
EQUATION:
Melissas + Christina's = $122
n + 3n+22 = 122
4n=100
n=$25 (amount Melissa has)
3n+22= 75+22= #97 (amount Christina has)
Cheers,
stan H.